# Detecting and removing bloated dependencies in CommonJS packages

### v2

@@ 1 Introduction
Such bloat issues also affect code written in JavaScript. Its dominant usage is facilitated by the Node Package Manager (npm 111https://www.npmjs.com), which provides reusable external packages (i.e., dependencies), but also contributes to code bloat. The portability of JavaScript and the ease of integrating new npm dependencies encourages fine-grained reuse, often resulting in the inclusion of small-sized dependencies, commonly referred to as “trivial packages” (e.g., is-odd222https://github.com/i-voted-for-trump/is-odd). While this fine-grained modularity promotes reusability, it also amplifies the prevalence of dependency bloat. Even small JavaScript projects can have large dependency trees, leading to maintenance and security challenges (Abdalkareem et al., 2020). Current research focuses primarily on mitigating code bloat at the function and file levels (Koishybayev and Kapravelos, 2020; Turcotte et al., 2022; Vázquez et al., 2019; Rollup, 2023). However, such fine-grained analyses do not support the total removal of unnecessary dependencies from the dependency specification file. Consequently, these dependencies stay in the dependency tree and still need to be maintained and evolved even though they are not used in the project.
@@
a dataset of 91 open-source CommonJS packages with strong test suites, accessible on Zenodo44410.5281/zenodo.14673020.
@@ 2.1 Structure of npm Packaging System
Developers modify only direct dependencies in the package.json file, without directly interacting with indirect dependencies listed in package-lock.json. In environments such as continuous integration setups, developers can ensure an exact installation of all dependencies with the corresponding versions specified in the package-lock.json file, by using the npm ci command. This guarantees that the package is built with the same dependencies every time, leading to consistent builds.
@@ 2.2 CommonJS Module System and the Need for Dynamic Analysis
(ii) As a function, require() can accept variable parameters, (iii) which may change from one code execution to the next It is common for developers to pass variables as parameters to the require() function. In LABEL:lst:require_in_runtime, this is illustrated by a plugin scenario, where the pluginName is computed in other parts of the code. The require() call itself is part of a function that can be invoked repeatedly, potentially loading different dependency packages with each execution.
@@ 3 Automatic Analysis of Bloated Dependencies in CommonJS Packages
A dependency is classified as unaccessed if none of its modules within the dependency are accessed by the OS file system during the execution of the package.
@@
A dependency is considered bloated if its complete removal does not impede the successful execution of the package.
Note that unaccessed dependencies differ from bloated dependencies. An unaccessed dependency is identified as not used by the OS file system during execution, however, it can be classified as bloated only if it can be entirely removed without causing any runtime failures.
@@ 3.2 Trace-based Bloat Detection
1.
Execution and tracing The process begins by executing the package using the input workload. During execution, our approach monitors all interactions between the OS file system and the running processes. This monitoring generates a detailed log of all file operations, including file paths across various formats, as illustrated in LABEL:lst:rec-executions. Only modules actively accessed by the OS file system during execution are captured, while those not involved in the execution are excluded.
2.
Filtering accessed modules We then filter the file paths to isolate those relevant to dependencies. Specifically, we target modules located in the dependency folder (package/node_modules) with extensions .js or .json,. This step generates a refined list of accessed modules, represented by path. For instance, line of LABEL:lst:rec-executions indicates that module airtap/node_modules/airtap-default/index.js is accessed during execution.
3.
Mapping accessed modules to dependencies Each accessed file path is then analyzed to determine its corresponding dependency. This step leverages the npm dependency installation mechanism, where each dependency is installed in a folder named after the dependency itself, as explained in Section 2.1. For example, in line of LABEL:lst:rec-files, the module airtap/node_modules/airtap-default/index.js maps to the dependency airtap-default. At this stage, we generate a list of accessed dependencies, capturing all dependencies with at least one module accessed during runtime.
4.
Identifying unaccessed dependencies Finally, our approach identifies unaccessed dependencies by comparing the complete set of runtime dependencies, with the list of accessed dependencies generated in the previous step. Dependencies that are not found in the accessed list are classified as unaccessed and considered candidates for debloating in the next phase.
@@ 3.3 Dependencies Debloating
1.
Removing unaccessed dependencies from metadata files The process begins by setting the initial set of bloated dependencies as the unaccessed dependencies identified in the trace-based analysis. These dependencies are located within the metadata files: the package.json file, which specifies direct dependencies, and the package-lock.json file, which records the entire dependency tree. For direct-only debloating, we modify the package.json file to remove declarations of direct unaccessed dependencies, ensuring they are no longer listed. This helps streamline the management of direct dependencies. For full-scale debloating, we modify the package-lock.json file to eliminate occurrences of both direct and indirect unaccessed dependencies. In both strategies, the removal of dependencies is conducted all at once, pruning the entire dependency tree in one step.
2.
Rebuilding the package Once the metadata files are updated, we use them to rebuild the package and verify the changes. In direct-only debloating, the rebuild uses the updated package.json file. This ensures that direct unaccessed dependencies are removed from their declaration file and excluded from future builds. In full-scale debloating, the rebuild uses the modified package-lock.json file. This prevents the installation of any unaccessed dependencies, including indirect ones, effectively pruning the dependency tree.
3.
Validating functionality After rebuilding, we execute the same inputted workload to verify that the package remains functional. This step ensures that the removal of dependencies does not compromise the behavior of the package. If functionality issues arise during validation, we restore the mistakenly removed dependencies and adjust the list of bloated dependencies to exclude them. The validation process is repeated until the debloating preserves the package’s original functionality.
@@ 3.4 Implementation
During the debloating phase, to restore dependencies, we copy the original package source code and the generated package-lock.json, then we use the official npm command npm ci to strictly install each dependency exactly as specified in the package-lock.json file (npm Docs, 2024). To produce debloated versions of package.json and package-lock.json, we leverage the Python built-in json library to parse the original files, identify bloated dependencies, and remove each occurrence of these dependencies from the dependencies of other packages. This involves removing direct bloated dependencies from the package.json, and both direct and indirect dependencies from package-lock.json.
@@ 4.2 Dataset Collection
To the best of our knowledge, our dataset is the first curated dataset of well-tested npm packages that use the CommonJS module system. We make this dataset publicly available on Zenodo: 10.5281/zenodo.14673020.
@@
Figure 2 provides an overview of our dataset, in which packages are ordered by lines of code (LoC) in ascending order. The upper part illustrates the number of runtime direct and indirect dependencies for each package. The number of indirect dependencies is counted based on unique installations, meaning that we count each instance wherever a dependency is installed multiple times. Across our dataset of 91 packages, there are a total of 50,488 runtime dependencies, including 869 direct and 49,619 indirect ones. The number of direct dependencies ranges from 1 (in kuzzle-plugin-mqtt) to 69 (in cedric05-cli), with a median of 7, while the number of indirect dependencies ranges from 6 (in caviar) to 1,521 (in easywebpack-js), with a median of 462. Most dependencies are not declared directly by developers but are instead pulled in indirectly through other packages, which aligns with previous studies (Soto-Valero et al., 2021b).
@@
In the lower part of Figure 2, we present the number of lines of code (LoC) in each package and the number of lines of code in their dependencies. The LoC count is obtained using cloc (AlDanial, 2023) and excludes JavaScript code comments. In total, the runtime dependencies across the 91 packages include more than 34 million lines of code. The LoC in packages ranges from 34 (in mkroosevelt) to 38,716 (in easywebpack-js), with a median of 410. The minimum number of LoC in the runtime dependencies is 1,771 (in caviar), while the maximum is more than 2 million (in easywebpack-vue). This observation is consistent with previous studies indicating that most of the code in a package comes from its dependencies rather than the original source code (Koishybayev and Kapravelos, 2020).
@@
We see from Figure 2 that neither the number of direct dependencies, nor the number of indirect dependencies, nor the LoC count in the dependencies are correlated with the package size. Notably, our dataset reveals that smaller packages are highly likely to introduce a significant number of dependencies, consistent with previous findings (Abdalkareem et al., 2020). For instance, the package brisky contains only 48 lines of its own source code, yet it imports 10 direct dependencies that lead to 433 indirect dependencies, resulting in a total of 298,840 lines of code from all dependencies. This underscores how small packages can accumulate extensive dependency trees, highlighting the potential for unnecessary dependencies that may not contribute to the package’s functionality.
@@
Although our minimum threshold for including a package in the dataset is 70% test coverage, the median coverage is 96.9%, with 32 packages achieving 100% coverage. This high level of test coverage makes our dataset well-suited for our evaluation, ensuring the relevance of our findings.
@@ 5 Experimental Results
To answer RQ1, we execute DepPrune across the 91 packages in our dataset, containing 10,268 test cases and 50,488 dependencies (direct and indirect). Our trace-based analysis determines that 135,095 modules (files), which belong to 24,972 dependencies, are accessed at runtime. Conversely, 25,566 dependencies comprise only modules that are never accessed at runtime and are therefore identified as unaccessed dependencies, as explained in Section 3.3. Overall, DepPrune detects that 50.6% (25,566/50,488) of all dependencies in our dataset are bloated.
@@
Although our focus is on removing direct bloated dependencies, we observe that indirect bloated dependencies can also exist independently of direct bloated dependencies, often introduced through non-bloated direct dependencies. For example, packages like generator-node-lib contain 798 indirect bloated dependencies despite having no direct bloated dependencies. Similarly, the package easywebpack-js (at row 46 in Table 2) includes 1,521 dependencies in its tree, with only one direct bloated dependency that does not contribute to the removal of any indirect dependencies. Among its indirect dependencies, 80.7% (1,227/1,521) are bloated and stem from its single non-bloated direct dependency (@easy-team/easywebpack). These findings demonstrate that indirect bloated dependencies can propagate through non-bloated direct dependencies, independent of direct bloat, highlighting the challenge of slimming dependency trees.
@@ 5.3.1 Comparison with Dynamic Analysis
We see that a significant proportion ( or 43.3%) of the dependencies identified by the coverage-based analysis are misclassified as bloated. These dependencies include cases where modules are loaded but are determined to be non-executed, causing them to be marked as unaccessed. To investigate the underlying reasons for these misclassifications generated by coverage-based analysis, we conduct a manual analysis of a random selection of 100 dependencies, using a 95% confidence level for our sampling.
@@ 5.4 Performance
DepPrune successfully executes and analyzes all 91 packages in the dataset, using their test suites as workloads. All identified bloated dependencies can be safely removed while preserving the packages’ functionality during evaluation. The additional time required by DepPrune, measured as a multiple of the original workload execution time, varies from 0.2x to 27.4x, with a median of 1.4x. For example in the package fastify-axios, the original workload takes 3.5s while running DepPrune on this package takes 4.2s, resulting in the additional 0.7s (0.2x) of the original workload. In addition, 22 packages have this overhead lower than 1x, demonstrating minimal additional time required for our analysis; while only 3 packages have the overhead higher than 10x. Notably, all packages with more than 100 test cases have an overhead lower than 2x. For example, the package cedric05-cli, which has 4,396 test cases, has the overhead only 1.04x. These results highlight that the overhead introduced by DepPrune is generally low, making it a practical and efficient solution for dependency analysis with manageable runtime overhead.
@@ 7.1 JavaScript Debloating
Our work differs from the above studies in several aspects. We conduct dynamic analysis on CommonJS packages, which include dynamically loaded modules mentioned in previous research (Koishybayev and Kapravelos, 2020; Ye et al., 2021). We evaluate our approach on a much larger dataset, using packages with high test coverage (median 96.9%). Our focus is removing entire dependencies from the dependency tree of a package, leaving the original package source code unchanged, a novel aspect not previously explored in JavaScript. Our approach can aid developers in reducing the maintenance costs associated with bloated dependencies, such as notification fatigue.

### v3

@@ 1 Introduction
Such bloat issues also affect code written in JavaScript. Its dominant usage is facilitated by the Node Package Manager (npm 111https://www.npmjs.com), which provides reusable external packages (i.e., dependencies), but also contributes to code bloat. The portability of JavaScript and the ease of integrating new npm dependencies encourages fine-grained reuse, often resulting in the inclusion of small-sized dependencies, commonly referred to as “trivial packages” (e.g., is-odd222https://github.com/i-voted-for-trump/is-odd) (Abdalkareem et al., 2017; Chowdhury et al., 2021). While this fine-grained modularity promotes reusability, it also amplifies the prevalence of dependency bloat. Even small JavaScript projects can have large dependency trees, which suffer from dependency overhead, leading to maintenance and security challenges (Abdalkareem et al., 2020). For example, the 11-lines left-pad package, is reported to pose a significant threat to the software supply chain (Open Source Security Foundation, 2023). Current research focuses primarily on mitigating code bloat at the function and file levels (Koishybayev and Kapravelos, 2020; Turcotte et al., 2022; Vázquez et al., 2019; Rollup, 2023). However, such fine-grained analyses do not support the total removal of unnecessary dependencies from the dependency specification file. Consequently, these dependencies stay in the dependency tree and still need to be maintained and evolved even though they are not used in the project.
@@
a dataset of 91 open-source CommonJS packages with strong test suites, accessible on Zenodo44410.5281/zenodo.15090140.
@@ 2.2 CommonJS Module System and the Need for Dynamic Analysis
(ii) As a function, require() can accept variable parameters, (iii) which may change from one code execution to the next. It is common for developers to pass variables as parameters to the require() function. In LABEL:lst:require_in_runtime, this is illustrated by a plugin scenario, where the pluginName is computed in other parts of the code. The require() call itself is part of a function that can be invoked repeatedly, potentially loading different dependency packages with each execution.
@@ 3 Automatic Analysis of Bloated Dependencies in CommonJS Packages
A dependency is classified as unaccessed if none of its modules are accessed during the execution of the package.
It is important to note that the classification of a dependency as unaccessed depends on the type of analysis used. For example, unaccessed dependencies can be detected by instrumenting individual functions within a dependency and monitoring whether those functions are invoked at runtime (Vázquez et al., 2019). Alternatively, they can be determined by analyzing execution reports, such as coverage reports, to identify dependencies that are not exercised during runtime (Turcotte et al., 2022). In this work, we develop a technique based on monitoring the file system to determine whether a dependency is accessed or not.
@@
An unacccessed dependency is considered bloated if its complete removal from the dependency tree does not impede the successful execution of the package.
We note here, that the key difference between a bloated and an unaccessed dependency is that a dependency can only be confirmed as bloated if it is uninstalled from the package and the package continues to function correctly after re-execution. Due to the highly dynamic and asynchronous nature of JavaScript, it is challenging for program analysis techniques to precisely and completely capture the dependencies accessed at runtime. For example, if the analysis relies on test coverage data (Section 2.3), asynchronous functions that execute after the coverage report is generated may be missed. As a result, the analysis may incorrectly classify a required dependency as unaccessed. In such cases, removing the misclassified dependency can break the program, indicating that it is not bloated despite being labeled unaccessed.
@@ 3.2 Trace-based Bloat Detection
1.Execution and tracing The process begins by executing the package using the input workload. During execution, our approach monitors all interactions between the OS file system and the running processes. This monitoring generates a detailed log of all file operations, including file paths across various formats, as illustrated in LABEL:lst:rec-executions. Only modules actively accessed by the OS file system during execution are captured, while those not involved in the execution are excluded.
2.Filtering accessed modules We then filter the file paths to isolate those relevant to dependencies.
Specifically, we target modules located in the dependency folder (package/node_modules) with extensions .js or .json.
This step generates a refined list of accessed modules, represented by path.
For example, line of LABEL:lst:rec-executions illustrates that the module airtap/node_modules
/airtap-default/index.js is accessed during execution.
3.Mapping accessed modules to dependencies Each accessed file path is then analyzed to determine its corresponding dependency. This step leverages the npm dependency installation mechanism, where each dependency is installed in a folder named after the dependency itself, as explained in Section 2.1. For example, line of LABEL:lst:rec-files illustrates that the module airtap/node_modules/airtap-default/index.js maps to the dependency airtap-default. At this stage, we generate a list of accessed dependencies, capturing all dependencies with at least one module accessed during runtime.
4.Identifying unaccessed dependencies Finally, our approach identifies unaccessed dependencies by comparing the complete set of runtime dependencies, with the list of accessed dependencies generated in the previous step. Dependencies that are not found in the accessed list are classified as unaccessed and considered candidates for debloating in the next phase.
@@ 3.3 Dependencies Debloating
1.Removing unaccessed dependencies from metadata files This process begins by setting the initial set of bloated dependencies as the unaccessed dependencies identified in the trace-based analysis. These dependencies are located within the metadata files: the package.json file, which specifies direct dependencies, and the package-lock.json file, which records the entire dependency tree. For direct-only debloating, we modify the package.json file to remove declarations of direct unaccessed dependencies, ensuring they are no longer listed. This helps streamline the management of direct dependencies. For full-scale debloating, we modify the package-lock.json file to eliminate occurrences of both direct and indirect unaccessed dependencies. In both strategies, the removal of dependencies is conducted all at once, pruning the entire dependency tree in one step.
2.Rebuilding the package Once the metadata files are updated, we use them to rebuild the package and verify the changes. In direct-only debloating, the rebuild uses the updated package.json file. This ensures that direct unaccessed dependencies are removed from their declaration file and excluded from future builds. In full-scale debloating, the rebuild uses the modified package-lock.json file. This prevents the installation of any unaccessed dependencies, including indirect ones, effectively pruning the dependency tree.
3.Validating functionality After rebuilding, we execute the same inputted workload to verify that the package remains functional. This step ensures that the removal of dependencies does not compromise the behavior of the package. If functionality issues arise during validation, we restore the mistakenly removed dependencies and adjust the list of bloated dependencies to exclude them. The validation process is repeated until the debloating preserves the package’s original functionality.
@@ 3.4 Implementation
During the debloating phase, we ensure that dependency resolution remains reproducible by copying the original package source code and the generated package-lock.json, then using the official npm command npm ci to strictly install each dependency exactly as specified in the package-lock.json file (npm Docs, 2024). To produce debloated versions of package.json and package-lock.json, we leverage the Python built-in json library to parse the original files, identify bloated dependencies, and remove each occurrence of these dependencies from the dependencies of other packages. This involves removing direct bloated dependencies from the package.json, and both direct and indirect bloated dependencies from package-lock.json.
@@ 4.2 Dataset Collection
To the best of our knowledge, our dataset is the first curated dataset of well-tested npm packages that use the CommonJS module system. We make this dataset publicly available on Zenodo: 10.5281/zenodo.15090140.
@@
Figure 2 provides an overview of our dataset, in which packages are ordered by lines of code (LoC), arranged in ascending order. The upper part illustrates the number of direct and indirect runtime dependencies for each package. The number of indirect dependencies is counted based on unique installations, meaning that we count each instance wherever a dependency is installed multiple times. Across our dataset of 91 packages, there are a total of 50,488 runtime dependencies, including 869 direct and 49,619 indirect ones. The number of direct dependencies ranges from 1 (in kuzzle-plugin-mqtt) to 69 (in cedric05-cli), with a median of 7, while the number of indirect dependencies ranges from 6 (in caviar) to 1,521 (in easywebpack-js), with a median of 462. Most dependencies are not declared directly by developers but are instead pulled in indirectly through other packages, which aligns with previous studies (Soto-Valero et al., 2021b).
@@
In the lower part of Figure 2, we present the lines of code (LoC) in .js or .ts files in each package and the lines of code in their dependencies. The LoC count is obtained using cloc (AlDanial, 2023) and excludes JavaScript code comments. In total, the runtime dependencies across the 91 packages include more than 34 million lines of code. The LoC in packages ranges from 34 (in mkroosevelt) to 38,716 (in easywebpack-js), with a median of 410. The minimum number of LoC in the runtime dependencies is 1,771 (in caviar), while the maximum is more than 2 million (in easywebpack-vue). This observation is consistent with previous studies indicating that most of the code in a package comes from its dependencies rather than the original source code (Koishybayev and Kapravelos, 2020).
@@
From Figure 2, we see that neither the number of direct dependencies, the number of indirect dependencies, nor the LoC count in the dependencies are correlated with the package size. Notably, our dataset reveals that smaller packages are highly likely to introduce a significant number of dependencies, consistent with previous findings (Abdalkareem et al., 2020). For instance, the package brisky contains only 48 lines of its own source code, yet it imports 10 direct dependencies that lead to 433 indirect dependencies, resulting in a total of 298,840 lines of code from all dependencies. This underscores how small packages can accumulate extensive dependency trees, highlighting the potential for unnecessary dependencies that may not contribute to the package’s functionality. Project size alone is therefore not a reliable indicator of project complexity.
@@
Although our minimum threshold for including a package in the dataset is 70% test coverage, the median coverage is 96.9%. Our dataset includes 32 packages with 100% coverage, 36 packages with 90–99.9% coverage, 15 packages with 80–89.9% coverage, and 8 packages with 70–79.9% coverage. This high level of test coverage makes our dataset well-suited for our evaluation, ensuring the relevance of our findings.
@@ 5 Experimental Results
To answer RQ1, we execute DepPrune across the 91 packages in our dataset, containing 10,268 test cases and 50,488 dependencies (direct and indirect). Our trace-based analysis determines that 135,095 modules (files), which belong to 24,972 dependencies, are accessed at runtime. Conversely, 25,566 dependencies comprise only modules that are never accessed at runtime and are therefore identified as unaccessed dependencies, as explained in Section 3.3. Overall, DepPrune detects that 50.6% (25,566/ 50,488) of all dependencies in our dataset are bloated, with a 100% of the test cases in all packages still passing after the their removal.
@@
Although this question focuses on removing direct bloated dependencies, we observe that indirect bloated dependencies can also exist independently of direct bloated dependencies, often introduced through non-bloated direct dependencies. These dependencies account for 83.1% (21,159 out of 25,446) of all indirect bloated dependencies, indicating that the majority of indirect bloat are not introduced through direct bloated dependencies, but instead originate from non-bloated direct dependencies. For example, the package generator-node-lib contains 798 indirect bloated dependencies despite having no direct bloated dependencies. Similarly, easywebpack-js (at row 46 in Table 2) includes 1,521 dependencies in its tree, with only one direct bloated dependency that does not contribute to the removal of any indirect dependencies. Among its indirect dependencies, 80.7% (1,227/1,521) are bloated and stem from its single non-bloated direct dependency. These findings demonstrate that indirect bloated dependencies can propagate through non-bloated direct dependencies, independent of direct bloat, highlighting the challenge of slimming dependency trees.
@@ 5.3.1 Comparison with Dynamic Analysis
We see that a significant proportion ( or 43.3%) of the dependencies identified by the coverage-based analysis are misclassified as bloated. These dependencies include cases where modules are loaded but are determined to be non-executed, causing them to be marked as unaccessed. To investigate the underlying reasons for these misclassifications generated by coverage-based analysis, we conduct a manual analysis of a random selection of 100 dependencies, using a 95% confidence level for our sampling. The margin of error for this sample is 9.77%.
@@ 5.4 Performance
DepPrune successfully executes and analyzes all 91 packages in the dataset, using their test suites as workloads. All identified bloated dependencies can be safely removed while preserving the packages’ functionality during evaluation. The additional time required by DepPrune, measured as a multiple of the original workload execution time, varies from 0.2x to 27.4x, with a median of 1.4x. For example, in the package fastify-axios, the original workload takes 3.5s while running DepPrune on this package takes 4.2s, resulting in the additional 0.7s (0.2x) of the original workload. In addition, 22 packages have this overhead lower than 1x, demonstrating minimal additional time required for our analysis. While 3 packages have an overhead higher than 10x, this occurs primarily in packages with a small number of test cases, where the original test execution time is very short, making the additional analysis time appear relatively large. Notably, all packages with more than 100 test cases have an overhead lower than 2x. For example, the package cedric05-cli, which has 4,396 test cases, has an overhead of only 1.04x. These results highlight that the overhead introduced by DepPrune is generally low, making it a practical and efficient solution for dependency analysis with manageable runtime overhead.
5.5 DepPrune in the Development Lifecycle
DepPrune debloats both direct and indirect dependencies. To do so, our tool automatically rewrites package.json and package-lock.json, enabling developers to build and install a package with no bloated dependencies. While package-lock.json is typically considered transient and may be regenerated during development, it plays a critical role in shaping the installed dependency tree. To prevent bloated indirect dependencies from reappearing, we recommend running DepPrune before each release. This ensures that the deployed environment includes only the necessary dependencies, resulting in a leaner dependency tree, without requiring changes from the npm package manager. This is particularly relevant in workflows that use bundlers (e.g., Webpack131313Webpack: a bundler for modern JavaScript applications), where the dependency tree directly impacts bundle size.
@@ 7.1 JavaScript Debloating
Our work differs from the above studies in several aspects. We conduct dynamic analysis on CommonJS packages, which include dynamically loaded modules mentioned in previous research (Koishybayev and Kapravelos, 2020; Ye et al., 2021). We evaluate our approach on a much larger dataset, using packages with high test coverage (median 96.9%). Our focus is on removing entire dependencies from the dependency tree of a package, leaving the original package source code unchanged, a novel aspect not previously explored in JavaScript. Our approach can aid developers in reducing the maintenance costs associated with bloated dependencies, such as notification fatigue.